# Automatic Input Feature Relevance via Spectral Neural Networks

### v2

@@ 4 Discussion
We have here elaborated on a procedure for feature relevance extraction, which is founded on solid mathematical grounds. The interest of the proposed method resides in that the user is not asked to decide a priori on what it is relevant or not. The assessment on the actual relevance of the provided input feature follows the evaluation of a scalar indicator (one for each of the features) that self consistently gauges the impact of the supplied information in the decision making problem. These parameters are mathematically traced back to the eigenvalues of the transfer operator linking the first two layers. The method is hence capable of self-adjusting to the specific problem under inspection, with no need for ad hoc (possibly biased) input from external operators.
Based on the above reported analysis carried out for both mock and real data, we can positively conclude on the adequacy of the proposed features’ detection algorithm.The eigenvalues associated to the input nodes of the collection provide indeed a sensible ranking of the input features.
By working with a relatively simple dataset, we could elaborate on a sound interpretation, drawn a posteriori, of the process that underlying decision making. This includes to visually isolate the relevant patches of information that are selectively seen by the trained device in operation mode. For more complex datasets, the posterior interpretation could be less transparent in terms of grasping the qualitative message delivered. As an obvious corollary of the analysis, the method could be also applied as a compression algorithm to store large data in a minimal memory kernel.
@@ 5 Methods
Spectral Formulation We provide here a few additional details on the spectral approach to the training [22]. As mentioned in the main body of the paper we consider layer $l-1$ and layer $l$ , respectively made of $N$ and $M$ nodes. The state of the system on layer $l-1$ is encapsulated in vector $\textbf{x}^{l-1}=\left(x^{l-1}_{1},\ldots,x^{l-1}_{N}\right)$. The state on layer $l$ is instead stored in vector $\textbf{x}^{l}=\left(x^{l}_{1},\ldots,x^{l}_{M}\right)$. The mapping is implemented as $\textbf{x}^{l}=f(W\textbf{x}^{l-1})$, where $W$ is a $M\times N$ matrix and $f$ The spectral recipe deals with the equivalent transformation $\textbf{z}^{l}=f(A\textbf{z}^{l-1})$, where $\textbf{z}^{l-1}=\textbf{x}^{l-1}=\left(x^{l-1}_{1},x^{l-1}_{2},\ldots x^{l-1}_{N},0,0,\ldots 0\right)$ and the last $M$ elements of vector $\textbf{z}^{l}$ refer to the actual activity displayed on layer $l$. An appropriate projection matrix should be formally considered at the end of each mapping between consecutive layers, which effectively implements the transfer of the activations from bottom to upper elements (while filling the bottom portion of the vector with zeros) of the next vector to be processed.
As anticipated in the main body of the paper $A=\Phi\Lambda\Phi^{-1}$ where $\Phi$ and $\Lambda$ are $M\times N$ matrices. The first identifies the eigenvectors of $A$, the latter the associated eigenvalues. More explicitly we assume:
@@
$\Lambda=\begin{pmatrix}\lambda_{1}&&&&\\ &\lambda_{2}&&&\\ &&\ddots&&\\ &&&&\lambda_{N+M}\\ \end{pmatrix}$ (5)
As already recalled it can be proven that $\Phi^{-1}=2\mathbb{I}-\Phi$ [26, 23]. By making use of this latter expression, it is immediate to show that:
@@
The part of $A$ responsible of the information transfer from the departure layer $l-1$ to the destination layer $l$ is the the bottom-left block $W$. The relation between the elements of $W$ and the spectral parameters (i.e., the non-zero elements of the matrix $\Lambda$ and the elements of the $\Psi$ block of the eigenvector matrix) can be analytically computed yielding
@@
Training procedure The optimization process for all three datasets discussed in the main text was conducted using the Adam optimizer [38]. The learning rates were set to 0.01 for the first two datasets, 0.0001 for MNIST and 0.001 for the stellar dataset. Sparse Categorical Crossentropy was employed as the loss function for the three classification tasks, while the mean absolute error was used as loss function for the metallicity prediction task. For all tasks under consideration, we applied regularization to both the eigenvalues and eigenvectors of the first layer. The regularization coefficients were set to $10^{-2}$ for the eigenvalue term and $10^{-4}$for the eigenvector term. The number of epochs was set to 1000 for the Gaussian dataset, 500 for the correlated Gaussian dataset, 5000 for MNIST and 1000 for the stellar spectra dataset. Using a large number of training epochs allows us to significantly reduce the regularization term associated with the eigenvalues in the loss function. This, in turn, enables the model to converge to a solution that relies on a minimal subset of input features while still correctly classifying the data. Such sparsity in the input improves the interpretability of the model. It is important to note that employing a large number of epochs is not a stringent requirement: even when operating with fewer training epochs or smaller regularization terms, the eigenvalues of the input layer remain valid indicators of node relevance. However, under these conditions, the model tends to distribute importance across a larger number of input components. While this still yields accurate classification, the interpretation of the internal representations becomes more diffuse and less localized.
@@
As empirical evidence of the above, we report the results for the MNIST dataset and the stellar dataset as obtained for a smaller number of training epochs and reduced regularization factors. By significantly reducing the number of epochs during training on the MNIST dataset, we find a bimodal distribution for the eigenvalues that refer to the input layer with a peak close to zero that is less pronounced as compared to that reported in Sec. (3.3), and a prominent bulk in the region of larger values (See left panel of Fig. 12). This indicates that a greater number of input components are being used to solve the task. By examining the spatial distribution of the eigenvalues, we observe that the relevant components are spread across nearly the entire image (See right panel of Fig. 12). In Figure (13) the results obtained for the stellar dataset are displayed: the eigenvalues associated to the input layer are plotted (black vertical lines) for a model trained with 100 epochs and a regularization coefficient for the eigenvalues set to $5\times 10^{-4}$.The test loss at the end of the training is essentially the same as that obtained for the case reported in Sec. (3.4) (not shown). However, the number of non-zero eigenvalues in the first layer is larger, although most of the peaks still lie in spectral regions known to be crucial for metallicity prediction. In other words the network detected another equivalent solution that makes use of a larger set of input components.
@@
We also evaluated the effect of different settings for other hyperparameters, such as the number of layers, the use of batch normalization, and the application of dropout. We did not observe any significant changes in the results under these variations. We would like to recall that our method is input-independent as it does not rely on the activation of the nodes to estimate their relevance (unlike methods such as LRP, for example). Additionally, the eigenvalues of the input layer are parameters directly tied to the input components and are unaffected by operations applied in the subsequent layers of the network. For these reasons, our method demonstrates strong robustness to architectural changes and to the use of techniques like batch normalization or dropout. Therefore, the hyperparameters were manually optimized from a restricted set of possible candidates by assessing the model’s performance on a validation set ($10$ of the training set).

### v3

@@ 4 Discussion
We have here elaborated on a procedure that allows one to easily extract input feature relevance in a neural network model and to force solutions relying on sparse representations. The procedure is input-independent and it is founded on solid mathematical grounds. The assessment on the actual relevance of the provided input feature follows the evaluation of a scalar indicator (one for each of the features) that self-consistently gauges the impact of the supplied information in the decision making problem. These parameters are mathematically traced back to the eigenvalues of the transfer operator linking the first two layers. The eigenvalues are trainable parameters and they are adjusted during the training to solve the task under consideration. During training, it is possible to regularize these parameters to enforce solutions making use of a minimal set of the input components.
Based on the above reported analysis carried out for both mock and real data, we can positively conclude on the adequacy of the proposed features’ detection algorithm. The eigenvalues associated to the input nodes of the collection provide indeed a sensible ranking of the input features. The relevance of input components obtained with the proposed method has been compared with that obtained working with different well-known methodologies showing good agreement.
By working with a relatively simple dataset, we could elaborate on a sound explanation, drawn a posteriori, of the process that underlying decision making. This includes to visually isolate the relevant patches of information that are selectively seen by the trained device in operation mode. For more complex datasets, the posterior explanation could be less transparent in terms of grasping the qualitative message delivered. As an obvious corollary of the analysis, the method could be also applied as a compression algorithm to store large data in a minimal memory kernel.
@@ 5 Methods
Spectral Formulation We provide here a few additional details on the spectral approach to the training [giambagli2021machine]. As mentioned in the main body of the paper we consider layer l1l-1 and layer $l$ , respectively made of $N$ and $M$ nodes. The state of the system on layer l1l-1 is encapsulated in vector $\textbf{x}^{l-1}=\left(x^{l-1}_{1},\ldots,x^{l-1}_{N}\right)$. The state on layer $l$ is instead stored in vector $\textbf{x}^{l}=\left(x^{l}_{1},\ldots,x^{l}_{M}\right)$. The mapping is implemented as $\textbf{x}^{l}=f(W\textbf{x}^{l-1})$, where $W$ is a $M\times N$ matrix and $f$ is a non-linear funtion. The spectral recipe deals with the equivalent transformation $\textbf{z}^{l}=f(A\textbf{z}^{l-1})$, where $\textbf{z}^{l-1}=\textbf{x}^{l-1}=\left(x^{l-1}_{1},x^{l-1}_{2},\ldots x^{l-1}_{N},0,0,\ldots 0\right)$ and the last $M$ elements of vector $\textbf{z}^{l}$ refer to the actual activity displayed on layer $l$. An appropriate projection matrix should be formally considered at the end of each mapping between consecutive layers, which effectively implements the transfer of the activations from bottom to upper elements (while filling the bottom portion of the vector with zeros) of the next vector to be processed.
As anticipated in the main body of the paper $A=\Phi\Lambda\Phi^{-1}$, where $\Phi$ and $\Lambda$ are $M\times N$ matrices. The first identifies the eigenvectors of $A$, the latter the associated eigenvalues. More explicitly we assume:
@@
$\Lambda=\begin{pmatrix}\lambda_{1}&&&&\\ &\lambda_{2}&&&\\ &&\ddots&&\\ &&&&\lambda_{N+M}\\ \end{pmatrix}$. (9)
As already recalled it can be proven that $\Phi^{-1}=2\mathbb{I}-\Phi$ [buffoni2022spectral, chicchi2021training]. By making use of this latter expression, it is immediate to show that:
@@
The part of $A$ responsible of the information transfer from the departure layer l1l-1 to the destination layer $l$ is the bottom-left block $W$. The relation between the elements of $W$ and the spectral parameters (i.e., the non-zero elements of the matrix $\Lambda$ and the elements of the $\Psi$ block of the eigenvector matrix) can be analytically computed yielding
@@
Regularization During training we employ an $L_{1}$ ora a $L_{2}$ regularization loss on the spectral parameters, $\phi$ and $\lambda$. This could also be understood as an implicit minimization of the Frobenius norm of the operator $W=\lambda\odot\phi$ as $W||_{F}=\sum_{i,j}\lambda_{i}^{2}\phi_{ij}^{2}$, therefore learning a low rank representation of the operator.
@@
Training procedure The optimization process for all four datasets discussed in the main text was conducted using the Adam optimizer [2015-kingma]. The learning rates were set to 0.01 for the first two datasets, 0.0001 for MNIST and 0.001 for the stellar dataset. Sparse Categorical Crossentropy was employed as the loss function for the three classification tasks, while the mean absolute error was used as loss function for the metallicity prediction task. For all tasks under consideration, we applied $L_{2}$ regularization to both the eigenvalues and eigenvectors of the first layer. We also tested $L_{1}$ regularization when it is specified in the text. The regularization coefficients were set to $10^{-2}$ for the eigenvalue term and $10^{-4}$ for the eigenvector term. The eigenvalues are initially set to one, following the initialization rule presented in [giambagli2024student]. The number of epochs was set to 1000 for the Gaussian dataset, 500 for the correlated Gaussian dataset, 5000 for MNIST and 1000 for the stellar spectra dataset. Using a large number of training epochs allows us to significantly reduce the regularization term associated with the eigenvalues in the loss function. This, in turn, enables the model to converge to a solution that relies on a minimal subset of input features while still correctly classifying the data. Such sparsity in the input improves the explainability of the model. It is important to note that employing a large number of epochs is not a stringent requirement: even when operating with fewer training epochs or smaller regularization terms, the eigenvalues of the input layer remain valid indicators of node relevance. However, under these conditions, the model tends to distribute importance across a larger number of input components. While this still yields accurate classification, the explanation of the process becomes more diffuse and less localized.
@@
As empirical evidence of the above, we report the results for the MNIST dataset and the stellar dataset as obtained for a smaller number of training epochs and reduced regularization factors. By significantly reducing the number of epochs during training on the MNIST dataset, we find a bimodal distribution for the eigenvalues that refer to the input layer with a peak close to zero that is less pronounced as compared to that reported in Sec. (2.3), and a prominent bulk in the region of larger values (See left panel of Fig. 15). This indicates that a greater number of input components are being used to solve the task. By examining the spatial distribution of the eigenvalues, we observe that the relevant components are spread across nearly the entire image (See right panel of Fig. 15). In Figure (16) the results obtained for the stellar dataset are displayed: the eigenvalues associated to the input layer are plotted (black vertical lines) for a model trained with 100 epochs and a regularization coefficient for the eigenvalues set to $5\times 10^{-4}$. The test loss at the end of the training is essentially the same as that obtained for the case reported in Sec. (2.4) (not shown). However, the number of non-zero eigenvalues in the first layer is larger, although most of the peaks still lie in spectral regions known to be crucial for metallicity prediction. In other words, the network detected another equivalent solution that makes use of a larger set of input components.
@@
We also evaluated the effect of different settings for other hyperparameters, such as the number of layers, the use of batch normalization, and the application of dropout. We did not observe any significant changes in the results under these variations. The eigenvalues of the input layer are parameters directly tied to the input components and are unaffected by operations applied in the subsequent layers of the network. For these reasons, our method demonstrates strong robustness to architectural changes and to the use of techniques like batch normalization or dropout. Therefore, the hyperparameters were manually optimized from a restricted set of possible candidates by assessing the model’s performance on a validation set ($10$ % of the training set).
Data availability
A public repository has been created containing the code used during the work: https://github.com/Lorenzochicchi/SpectralNNs_for_Input_Relevance. Specifically, the code for generating the synthetic datasets used, the code for training the models, and the code for plotting the images included in the work are provided in the form of notebooks.
@@ Acknowledgements
Appendix A Gaussian dataset with non-isotropic covariance
We briefly discuss here a simple generalization of the Gaussian dataset analyzed in the main body of the paper. If the data are characterized by a non-isotropic covariance matrix, the distance between the one-dimensional distributions shown in Figure (2) is no longer a good estimator of the relative relevance of each input component, since the information is generally distributed across components. As an example, let us consider a simple case of an input vector composed of only two components, where only one shows partial separation between the two classes, when examining each component separately. Therefore, by inspecting the distributions of the two components, in the spirit of what was done for the Gaussian dataset in the main body of the paper, one might be led to think that only the second component is relevant for solving the task (panel (a) of Figure (17)). However, this is true only if the two components are independent (panel (b) of Figure (17)). When the two components are correlated instead (panel (c) of Figure (17)), the model must take both into account to solve the task efficiently. The post-training eigenvalues associated with the two components, shown in the contour plots for both cases, correctly identify these two different situations. Obviously, this is still a very simple (linearly separable) example, and the results can be compared with other relevance measures obtained from alternative methods, such as the coefficients of a LASSO regressor, which lead to analogous results (data not shown).
Figure 17: In panel (a), the marginal distributions for components $x$ and $y$ are shown. This situation could, in principle, correspond either to the case of uncorrelated components (panel (b)) or to that of correlated components (panel (c)). In the first scenario, component $x$ is irrelevant for solving the task, and its corresponding eigenvalue is set to zero after training (see the inset in the right image of panel (b)). In the second scenario, component $x$ is essential for correctly separating the classes, and two comparable eigenvalues are associated with the two components.
Appendix B Nonlinearly separable dataset with spurious features
In this appendix, we evaluate the performance of the proposed method on a synthetic dataset that includes spurious input components. A classification task with two distinct classes is simulated. Operatively, each input vector in the dataset consists of ten components: only the first component is used to define the class label, while the others are merely correlated with the first one and are not directly involved in determining the class. Therefore, the underlying problem is essentially a one-dimensional classification task.
To work in a setting more suitable for neural networks, the generated classes are not linearly separable. Specifically, we randomly generate the elements of the first component from a uniform distribution between 0 and 1, and we divide this interval into four equal subintervals, assigning each alternately to one of the two classes. Then, we generate the remaining components to be proportional to the first one. Formally:
$\begin{split}x_{0}^{d}&\in[0,1],\\ x_{i}^{d}&=x_{0}\alpha_{i}\beta^{d}+\epsilon_{i}+\eta^{d},\\ y^{d}&=\begin{cases}0\;\;\text{if}\;\;x_{0}^{d}\in[0,0.25)\cup[0.5,0.75)\\ 1\;\;\text{if}\;\;x_{0}^{d}\in[0.25,0.5)\cup[0.75,1)\end{cases},\end{split}$ (12)
where the index $i$ refers to the components and the index $d$ refers to the data samples. The parameters $\alpha_{i}$ and $\epsilon_{i}$ are random values that apply a linear transformation to the first component, while $\beta^{d}$ and $\eta^{d}$ are additional random values that may vary across data samples. If $\beta^{d}=1$ and $\eta^{d}=0$ for all samples, the nine additional components are perfectly correlated with the first one, and the problem can be equivalently represented in the domain of any component. Conversely, if these two parameters vary among the data, the correlation between the additional components and the first one becomes weaker or stronger depending on their variability.
We define the correlation strength $s$ between the first component and the others as the average value of the off-diagonal elements of the covariance matrix of the data. The proposed method was tested on different realizations of the dataset while exploring a range of possible correlation strengths.
First of all, the method always identifies only one component as relevant, since the information contained in the other components is redundant and the regularization term penalizes solutions that rely on multiple inputs. Moreover, Figure 18 reports the larger eigenvalue obtained from 10 different experiments and 10 distinct correlation strengths. Even at very high correlation strengths, the first component is correctly identified as the relevant feature. It is only when $s$ approaches 1 (perfect correlation) that the method randomly selects one of the equivalent components to solve the task.
Figure 18: The color represents the larger eigenvalue obtained in each of the 10 different experiments conducted, across 10 different correlation strengths.
Appendix C Methods comparison on non-isotropic Gaussian dataset
We report here a quantitative comparison between our method and three of the methods described in Section (3) for a Gaussian dataset. Specifically, a Gaussian dataset consisting of two classes and 100 input features was generated, following the approach presented in Section (2.1). The distributions associated with the two classes are partially separated in the first 20 features, which are therefore intrinsically relevant. The next 20 features show no separation between classes but were generated to exhibit a certain degree of correlation with some of the first 20 features and, as explained in Appendix (A), they are consequently relevant for solving the task. Finally, the remaining 60 features are spurious and contain no useful information.
We tested the performance of the spectral method, the saliency map, the integrated gradients, and the SHAP method on this class of datasets. In particular, we measured both the accuracy in detecting relevant and non-relevant components and the computation time of each method. The simplicity of this linearly separable dataset allows us to use a LASSO logistic regression model as a benchmark for the number of relevant components. As expected, the LASSO logistic regression identifies the first 40 components as the relevant ones.
Since all methods, except for the spectral method, are input-dependent, the global relevance estimate must be obtained by averaging results over multiple examples. Therefore, we report the outcomes obtained for different numbers of samples used to compute the global relevance. Figure (19) shows the results for 10 independent experiments in two scenarios. In the first one (panels (a) and (c)), all four methods are evaluated on the same model: a spectral neural network trained with an $L_{1}$ regularization on both eigenvalues and eigenvectors. However, because the $L_{1}$ regularization on eigenvalues is part of the proposed method itself, we also compare the results obtained with the spectral method with what obtained when the other methods are applied to a standard MLP with an $L_{1}$ regularization on the weights of the first layer (panels (b) and (d)).
In both scenarios, the spectral method proves to be the fastest, with a computation time several orders of magnitude smaller than that of the other methods. The accuracy scores are comparable across all the methods in the first scenario when an adequate number of examples are used to compute the global relevance. Conversely, the accuracy of the other methods when they are tested on standard MLP is considerably lower than that of the spectral method.
Figure 19: Panels (a) and (b) show the computation time required to estimate the global relevance of input components for the Gaussian dataset using the four analyzed methods: spectral method, saliency map, SHAP, and integrated gradients. Panels (c) and (d) show the accuracy of the same methods in identifying the relevant components.The results in panels (a) and (c) correspond to the case in which all methods are evaluated on the same spectral neural network (trained with an $L_{1}$ regularization on eigenvalues and eigenvectors). Panels (b) and (d) instead refer to the case where the standard methods (saliency map, SHAP, and integrated gradients) are evaluated on a multilayer perceptron (MLP) with the same architecture as the spectral network. During MLP training, an $L_{1}$ regularization term is applied to the weights of the first layer to promote sparse solutions.
Appendix D Algorithmic complexity of different Methods
We compare here the computational complexity of the methods presented in Section (3) with that of the spectral method. In the spectral method, the relevance is computed by multiplying the vector of eigenvalues from the first layer by the norm of the corresponding eigenvector entries. The algorithmic complexity of this operation is $O(n_{1}\times n_{2})$, where $n_{1}$ and $n_{2}$ are the numbers of nodes in the first and second layers, respectively. The algorithmic complexity of the saliency map is on the order of a single forward (or backward) pass [simonyan2013deep, ancona2018towards], denoted as $C_{f}$, which is larger than $n_{1}\times n_{2}$ for a multilayer perceptron (MLP). The LRP method exhibits the same complexity [montavon2018methods]. The Integrated Gradients method introduces an additional cost due to the $m$ steps required to approximate the path integral [sundararajan2017axiomatic, kapishnikov2021how]. Finally, the computational cost of SHAP is $O(2^{n_{1}}C_{f})$ [lundberg2017unified].
Recall that, except for the spectral method, all the other methods are designed to provide input-dependent explanations. Therefore, to obtain a global relevance estimate for the input components, it is necessary to average the results over $E$ examples. Table (1) summarizes the computational complexity of all the methods considered.
Table 1: Computational complexity of local and global relevance computation for various explanation methods.
Method Local Complexity Global Complexity (E examples)
Saliency Maps $O(C_{f})$ $O(E\cdot C_{f})$
Layer-wise Relevance Propagation (LRP) $O(C_{f})$ $O(E\cdot C_{f})$
Integrated Gradients $O(m\cdot C_{f})$ $O(E\cdot m\cdot C_{f})$
SHAP (exact) $O(2^{n_{1}}\cdot C_{f})$ $O(E\cdot 2^{n_{1}}\cdot C_{f})$
Spectral Method - $O(\cdot n_{1}\cdot n_{2})$
Appendix E Effect of training set size and regularization strength on method performance
We analyze here the performance of the proposed method on the first dataset introduced in the manuscript (independent Gaussian dataset) for different training set sizes and various values of the regularization factor applied to the eigenvalues. Figure (20) reports the post-training eigenvalues as a function of the relevance defined in Section (2.1) for different training set sizes (panel (a)) and different $L_{2}$ regularization strengths (panel (b)). We also compute the correlation between the eigenvalues and the relevance for all the considered cases (panels (c) and (d)).
The correlation decreases when the training set becomes sufficiently small, likely due to sampling fluctuations during dataset generation, which can lead to situations where the actual relevance of the components differs from the theoretical one. This interpretation is supported by the large standard deviations observed in this regime.
Regarding the effect of the regularization factor, the correlation remains high for all tested values, indicating that the eigenvalues are organized consistently with the relevance of the corresponding nodes and that their relative ordering is preserved. However, the absolute magnitudes of the eigenvalues are influenced by the strength of the regularization, as can be observed in panel (b).
Figure 20: Panels (a) and (b) show the relationship between the eigenvalues and the relevance for different training set sizes and regularization factors, respectively. Panels (c) and (d) show the correlation between the eigenvalues and the relevance scores for the same cases.